In Rectangle ABCD angle cbd=35, diagonals intersect at o .how to find angle AOB?

Answer by stanbon(48535)

In Rectangle ABCD angle cbd=35, diagonals intersect at o .how to find angle AOB?
------------
Draw the figure so that angle cbd = 35 degrees
Then angle cba = 55 because it is the complement of cbd.
And angle dac = 55 because of isosceles triangle CAO
------
Finally AOC = 180 - 2(55) = 180-110 = 70 degrees.
====================================================
Cheers,
Stan H.